Evaluate the following given in the p/q form
23 1/3%
71 3/7%
63 2/4%
17 3/4%
7 3//6%


Answer by stanbon(75887)

Evaluate the following given in the p/q form
23 1/3% = (3*23+1)/300 = 70/300 = 7/30
----------------------------------------------
71 3/7% = (7*71+3)/700 = 480/700 = 48/70 = 24/35
----------------------------------------------
63 2/4% = (4*63+2)/400 = 254/400 = 127/200
----------------------------------------------
Etc.
Cheers,
Stan H.
17 3/4%
7 3//6%